July 1-4

– Five stores provided ~ 64.5 hours of luminosity
– Tevatron sector E48 kicker caused problems
– NuMI dehumidifier shut off due to problems for weekend
– Power glitch tripped off many devices
– MI-12 service building high temperature caused MiniBooNE to turn off for seven hours
– A1 transfer line suffered a vacuum burst
– Linac RF station (LRF2) problems held off beam for an hour

*The integrated luminosity for the period from 6/27/11 to 7/4/11 was 59.43 inverse picobarns. NuMI reported receiving 5.12E18 protons on target during this same period.

Read the Current Accelerator Update
Read the Early Bird Report
View the Tevatron Luminosity Charts